Grade 12Physical Chemistry

What is the no. of atoms per unit cell (z) in a body-centred cubic structure???

aditya
5 Years agoGrade 12

1 Answer

Pawan Prajapati
5 Years ago
Answer:
Contribution by the atoms present at eight comers = 8 × 18 = 1
Contribution by the atoms present at centre = 1
Total number of atoms present in unit cell = 1 + 1 = 2

Thanks